what is the value of "a" in the equation a=xy2 when x=2 and y=2. would like individual steps for arriving at the solution. Thanks
Answer by stanbon(75887)

what is the value of "a" in the equation
a=xy^2 when x=2 and y=2
-----
a = 2*2^2
---
a = 2*4
---
a = 8
=======
Cheers,
Stan H.